OpenStudy (anonymous):

limit n tend to( n^2*e^-n)the sequences is converge or diverge

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \frac{x^2}{e^x}\]

OpenStudy (anonymous):

use Hopital rule twice

OpenStudy (anonymous):

can u solve it

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \frac{x^2}{e^x}=\lim_{x \rightarrow \infty} \frac{2x}{e^x}=\lim_{x \rightarrow \infty} \frac{2}{e^x}=\lim_{x \rightarrow \infty} \frac{2}{\infty}=0\]

OpenStudy (anonymous):

then the the sequences is converge

OpenStudy (anonymous):

thanks borther.. jazak ALLAH

OpenStudy (anonymous):

MARHABA

OpenStudy (anonymous):

BROTHER IN THE 3RD STEP ONLY 2/e^x where is x..? i mean 2x rather tthen x why.. can u explain it.

OpenStudy (anonymous):

oho brother again derivative..?

OpenStudy (anonymous):

plz bro ans this question

OpenStudy (turingtest):

please post each question separately thanks :)